Petra’s custom ornaments personalizes ornaments on the spot. She has a small kiosk at the mall, where customers walk up, select the ornament and specify the personalization. Petra can have up to three employees. An inexperienced employee takes 5 minutes to personalize each ornament is paid $12/hour and works 6 hour days each. She can hire an experienced personalizer who can produce the personalization in 2 minutes per ornament for $18/hour for a 6 hour day. If she has a request for 40 ornaments to be personalized each hour, calculate a) waiting time for each ornament, using the various combinations of employees, and b) identify what option gives the best utilization of employee time.
Answer #1

assuming the personalizer do not go for a break then

for inexperienced personalizer: total ornaments that can be processed in 1hr by 1 person = 60/5=12

for experienced personalizer: total ornaments that can be processed in 1hr by 1 person = 60/2=30

a. below is the average waiting time for each ornament under different scenarios. it is obtained by dividing the total ornaments that needs to be produced by total ornaments processed in 1 hr.

Total Ornaments processed in 1 hr Total Cost Incurred
Possible Options Experienced Inexperienced Total Employees Experienced Inexperienced Total Experienced Inexperienced Total Average waiting time
1 3 0 3 90 0 90 54 0 54 0.666666667
2 2 1 3 60 12 72 36 12 48 0.833333333
3 1 2 3 30 24 54 18 24 42 1.111111111
4 0 3 3 0 36 36 0 36 36 1.666666667

b. since 60 ornaments has to be personalised 2 experienced and 1 inexperienced employee combination is to be used for the same.

Total Ornaments processed in 1 hr Total Ornaments processed in 1 hr
Possible Options Experienced Inexperienced Total Employees Experienced Inexperienced Total Experienced Inexperienced Total
1 3 0 3 90 0 90 54 0 54
2 2 1 3 60 12 72 36 12 48
3 1 2 3 30 24 54 18 24 42
4 0 3 3 0 36 36 0 36 36